Opened 10 years ago

Closed 8 years ago

Last modified 2 years ago

#270 closed flight processing (fixed)

GB09/05, flight day 278b, Leighton moss

Reported by: mark1 Owned by: knpa
Priority: alpha 4 low Milestone: 2009 Data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by mark1)

Data location: ~arsf/arsf_data/2009/flight_data/uk/GB09_05-2009_278b_Leighton_moss

Data arrived from ARSF via SATA disk LETTER L on 7th October 2009.

Scientific objective: /users/rsg/arsf/arsf_data/2009/ARSF_Applications/GB09-05.pdf

Priority: alpha-4-low

PI: A. Blackburn

Sensors:

  • Eagle (26/02/2010)
  • Hawk (26/02/2010)
  • Leica LIDAR (26/02/2010)

Change History (27)

comment:1 Changed 10 years ago by mark1

Still awaiting logsheet. When it arrives, need to go through the eagle and hawk files and see why they have been put in 2 separate folders (278b and 278c) and move/rename as necessary. Can then create a DEM.

comment:2 Changed 10 years ago by mark1

logsheet arrived and unpacked

comment:3 Changed 10 years ago by mark1

Have split the project up into 2 separate ones (GB09/05 278b and 278c). Both contain duplicates of the same applanix and rinex data but have independent logsheets and raw eagle/hawk data. The Leica lidar data has been kept as one data set in this project (GB09/05 278b). If when it has been processed one can tell if it should be split (and into which project to put it) then do so. See ticket #279

comment:4 Changed 9 years ago by mggr

Re: #281 (Hawk frame rates are double what they should be), the following lines are affected and the .hdr file will need to be edited to halve the fps number:

uk/GB09_05-2009_278b_Leighton_moss/hawk/SWIR278b-09-1.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB09_05-2009_278b_Leighton_moss/hawk/SWIR278b-09-2.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB09_05-2009_278b_Leighton_moss/hawk/SWIR278b-09-3.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB09_05-2009_278b_Leighton_moss/hawk/SWIR278b-09-4.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB09_05-2009_278b_Leighton_moss/hawk/SWIR278b-09-5.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB09_05-2009_278b_Leighton_moss/hawk/SWIR278b-09-6.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB09_05-2009_278b_Leighton_moss/hawk/SWIR278b-09-7.hdr: predicted length half recorded length, frame rate doubling problem?
uk/GB09_05-2009_278b_Leighton_moss/hawk/SWIR278b-09-8.hdr: predicted length half recorded length, frame rate doubling problem?

comment:5 Changed 9 years ago by mark1

IPAS processing has been done.

comment:6 Changed 9 years ago by mark1

Started Hyperspectral processing

comment:7 Changed 9 years ago by mark1

None of the Hawk lines have specim sync messages or dark frames.

comment:8 Changed 9 years ago by mark1

  • Owner set to mark1
  • Status changed from new to assigned

comment:9 Changed 9 years ago by emca

Starting LIDAR processing.
Data in raw_laser relates to 278c, raw_laser for this project was bundled in with 278a.
Have moved the raw_laser files into the correct directories.

comment:10 Changed 9 years ago by emca

LIDAR processing complete. Ready for delivery check at:
~airborne/workspace/GB09_05-2009_278b_Leighton_moss/delivery/20100218/GB09_05

comment:11 Changed 9 years ago by mark1

Eagle LineSCT Offset
10
20
30.03
40.03
50.03
60
70
80
Hawk LineSCT Offset
1-4.7
2-4.6
3-4.4
4-5.3
5-5.05
6-5.24
7-4.45
85.2

comment:12 Changed 9 years ago by mark1

Hyperspectral delivery created at:

~airborne/workspace/GB09_05-2009_278b_Leighton_moss/delivery/20100223/GB09_05

comment:13 Changed 9 years ago by mark1

LIDAR delivery checked. All lines open up in grass and look good and data checked for missing records. Ready to deliver.

comment:14 Changed 9 years ago by knpa

  • Owner changed from mark1 to knpa
  • Status changed from assigned to accepted

comment:15 Changed 9 years ago by knpa

Hyperspectral delivery checked

All fine apart from:

  • Readme contents list; should be gap between last lev1 and misc listings
  • Cloud overflow on e3
  • Lots of underflows in H08
  • Need to mention blank band 232 in all hawk
  • Overflows on h03
  • Serious underflows on H06 - last 3 appear blank

comment:16 Changed 9 years ago by knpa

  • Owner changed from knpa to mark1
  • Status changed from accepted to assigned

comment:17 Changed 9 years ago by mark1

  • Description modified (diff)

Hyperspectral and LIDAR data delivered to PI on 26/02/2010

comment:18 Changed 9 years ago by mark1

rsyncing LIDAR and hyperspectral deliveries to ~arsf.

comment:19 Changed 9 years ago by mark1

  • Component changed from Processing: general to Archiving

Ready for archiving.

comment:20 Changed 8 years ago by mark1

File darkframes_from_SWIR278c-09-1.hdr appears to have wrong information for the raw file size. Have edited the hdr file (assumed number of lines was incorrect - changed from 30 to 15).

comment:21 Changed 8 years ago by knpa

  • Owner changed from mark1 to knpa

Beginning archiving.

comment:22 Changed 8 years ago by knpa

No webcam data.

comment:23 Changed 8 years ago by knpa

Created tarball, awaiting NEODC download.

comment:24 Changed 8 years ago by knpa

Backed up by NEODC.

Moved to archive disk.

comment:25 Changed 8 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Deleted workspace version.

Closing ticket.

comment:26 Changed 2 years ago by lah

Archiving

Tidied and ready for upload to NEODC.

comment:27 Changed 2 years ago by asm

Archiving

Re-processed hyperspectral data has been archived and everythin is ready to be downloaded from CEDA. Will close this ticket.

Note: See TracTickets for help on using tickets.